Home‎ > ‎Crew 169 Team Peru IV‎ > ‎

0513-DailySummaryReport

Crew 169 DAILY SUMMARY REPORT
SOL: (12)
Person filling out Report: XO
Summary Title: Sampling Phase IIIb and IIb.
Mission Status: Phase IIIb and IIb.

Sol Activity Summary: Today I’m sampling of the Phase IIIb and IIb. In the Phase IIIb I was sampling the Rover recovery of HSO, and on wheels of ATV. In the Phase IIb I was realizing a conventional sterilizing on the suit of two Crew members (HSO and Journalist) and the boots, after I sampling this points and comparing the presence of microbes.

Look Ahead Plan: The process consist in sampling these points and isolate on Petrifilm media and before incubated these media for 24 hours to 37±1°C.

Anomalies in work: No problems.
Weather: Phase Phase IIIb and IIb.
Crew Physical Status: Adequate.
EVA: Initial (2).
Reports to be file: Commander Report, XO Report, Engineering Report, HSO report.